OpenStudy (anonymous):

divide x^2+8x+15/4x^2+23x+15 divide x^2+6x+9/4x^2-5x-6

myininaya (myininaya):

\[\frac{x^2+8x+15}{4x^2+23x+15} \cdot \frac{4x^2-5x-6}{x^2+6x+9}\] \[\frac{(x+5)(x+3)}{4x^2+20x+3x+15} \cdot \frac{(x-6)(x+1)}{(x+3)(x+3)}\]

myininaya (myininaya):

now i basically have done this problem for you i left some factoring for you in the denominator of the first fraction(hint factor by grouping) after that you will cancel some stuff

OpenStudy (anonymous):

i m not getting any where

myininaya (myininaya):

on the part I left you?

myininaya (myininaya):

do you know how to factor by grouping?

myininaya (myininaya):

\[4x^2+20x+3x+15=4x(x+5)+3(x+5)\] can you do it from here?

myininaya (myininaya):

just factor 4x(x+5)+3(x+5)

myininaya (myininaya):

and you will be done with the factoring part then you can cancel stuff